how do you solve 2/3m+2=-3+3/2m?
Answer by nyc_function(2741)

2/3m+2 = -3+ 3/2m

To remove the fraction, multiply all terms on both sides of the fractional equation by the LCD 6m.

(2/3m)*6m + 2*6m = -3*6m + (3/2m)*6m

4 + 12m = -18m + 9

12m + 18m = 9 - 4

30m = 5

m = 5/30

m = 1/6